OpenStudy (anonymous):

a^14/3 x a^5/6= Use the law of exponents to simplify

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\]

OpenStudy (anonymous):

the exponents are fractions what do i do

OpenStudy (unklerhaukus):

to add these fractions \[\frac{14}3+\frac{5}{6}\] convert the first fraction to 6ths

OpenStudy (anonymous):

16/6+5/6

OpenStudy (unklerhaukus):

\[\frac{14}3=\frac{14}3\times1=\frac{14}3\times\frac22=\frac{14\times2}{3\times2}\]

OpenStudy (anonymous):

28/6+5/6

OpenStudy (anonymous):

33/6

OpenStudy (unklerhaukus):

Yes!

OpenStudy (unklerhaukus):

so what is your final answer?

OpenStudy (anonymous):

11/2

OpenStudy (unklerhaukus):

yeahs, so a^(14/3) * a^(5/6) =

OpenStudy (anonymous):

11/2

OpenStudy (unklerhaukus):

\[\large a^{8/3} \times a^{5/6}=a^{(14/3+5/6)}\\ \large\qquad\qquad=a^{(28/6+5/6)}\\\large\qquad\qquad=a^{(28+5)/6}\\\large\qquad\qquad=a^{33/6}\\\large\qquad\qquad=a^{11/6}\]